OpenStudy (anonymous):

You are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x) = -x^4 + x^3 + 7x^2 - x - 6. Describe to the CEO what the graph looks like. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0).

OpenStudy (mathmale):

Make certain you know by heart what the graphs of y=c, y=x, y=x^2, y=x^3 and y=x^4 look like, and what "end behavior" means in this context. Notice that the given function is a 4th order polynomial. Will its graph open up or down? Why? Where will this graph cross the x-axis? These points are your "zeros" or "x-intercepts" or "roots." Your break-even point will be one of these roots. If you can follow these directions, you'll be well on your way to solving this problem. If you can't, look up the terminology and/or ask specific questions here on OpenStudy.

OpenStudy (mathmale):

A warm welcome to OpenStudy!